OpenStudy (anonymous):

What are the asymptotes of the hyperbola given by the equation y^2/1- x^2/121 = 1?

OpenStudy (agent0smith):

You can use the same formula as the last question since it's another vertical hyperbola. This time h and k are zero, a^2=1 and b^2=121.

OpenStudy (agent0smith):

\[y=\pm \frac{ a }{ b }(x-h)+k \]

OpenStudy (abb0t):

hyperbola with vertical transverse axis. The standard form is: \[\frac{ (y-k)^2 }{ a^2 } - \frac{ (x-h)^2 }{ b^2 } = 1\] where (h,k) = (x,y) respectively. And are the coordinates of center. For your problem: center: (0,0) a = 1 b = 121 = 11 (it's 11 because it's b^2, hence, b = 11. You can't have b^2). Now, remember, asymptotes are straight lines that go thru the center (0,0). While your standard form of equation for a straight line is defined by: y = mx + b m = slope b = y-intercept Slopes of asymptotes are: \[\pm \frac{ a }{ b } = \frac{ 1 }{ 11 }\] hence: \[y = \pm \frac{ x }{ 11 } + b\] since your center is at (0,0), y-intercept = 0 b = 0 and your equation of asymptotes is: \[y = \pm \frac{ x }{ 11 }\]

OpenStudy (anonymous):

choices are: y=121x y=11x y= 1/11x y= 1/121x

OpenStudy (anonymous):

@abb0t which one is it?

OpenStudy (agent0smith):

which one does \[y = \pm \frac{ x }{ 11 } \] look like?

OpenStudy (abb0t):

\[y = \pm \frac{ x }{ 11 } = \pm \frac{ 1 }{ 11 }x\]